OpenStudy (goformit100):

If the HCF of 65 and 117 is expressible in the form 65m– 117, then the value of m is (A) 4 (B) 2 (C) 1 (D) 3

OpenStudy (goformit100):

@dan815

OpenStudy (anonymous):

prime factor of 65 =5*13 prime factor of 117 =3*3*13 --------------------------------------… H.C.F. of 65 and 117 = 13 according to question 65m -117 = 13 65m = 13+117 65m =130 m =130/65 =2 ANS.